# Clone Deterministic 3D Worlds

### v2

@@ 5.2 Temporal Contrastive Regularization
The output of the encoder, a sequence of latent vectors $\mathbf{z}\in\mathbb{R}^{B\times L\times\dots}$, is first passed through a projection layer (a linear layer) to produce embeddings $\mathbf{p}\in\mathbb{R}^{B\times L\times D}$ [^4] BB denotes the batch size, $L$ the sequence length, and $D$ the embedding dimension. (chen2020simple). These embeddings are then $L_{2}$-normalized to lie on the unit hypersphere: $\mathbf{p}^{\prime}=\frac{\mathbf{p}}{\|\mathbf{p}\|_{2}}$ (wang2020understanding). Our regularization objectives are applied to these normalized embeddings, $\mathbf{p}^{\prime}$.
@@
The idea of temporal slowness is that consecutive or nearby states in a trajectory should have similar latent representations, reflecting the intuition that the underlying state of the environment evolves gradually over time (wiskott2002slow). Our loss encourages all pairs of frames within the same trajectory’s context window to be close to one another on the hypersphere. This enforces that the entire trajectory segment is mapped to a compact and continuous path in the representation space, ensuring that the latent representation evolves slowly and smoothly over time. We formalize this by minimizing the average L2 distance between all pairs of embeddings within a trajectory:
@@ 6.1 Setup
State-of-the-art latent generative world models generally adopt a standard VAE as the backbone for representation learning. Our method does not modify the dynamics model itself. Therefore, when comparing with existing sota approaches, we pair our representation module with the same dynamics models they use. We select three sota latent generative world models: Standard Diffusion (SD) (alonso2024diffusionworldmodelingvisual), Video Diffusion Models (VD) (ho2022video), and Diffusion Forcing (DF)  (chen2024diffusion). A detailed description of training configurations is provided in Appendix B.
@@ 6.3 Qualitative Long-Horizon Results
To further probe the long-term stability, we conduct an extreme long-horizon generation task. For this experiment, we specifically select sota world model, Diffusion Forcing [^5] In the following discussion, we refer to the original Diffusion Forcing as VAE-WM, and GRWM with Diffusion Forcing as GRWM., and the M9x9-DET and MC-DET dataset. We use a sequence of randomly sampled actions to simulate an trajectory.
We evaluate our model on trajectories of varying lengths. To test long-horizon stability, we tasked the model with producing 10,000 consecutive frames from a single starting point. We visualize these long trajectories in Figure 4 and Figure 6. For the more complex MC-DET environment, we specifically evaluate a 100-frame rollout, as shown in Figure 5. Additional results under different starting conditions are provided in the Appendix A.
@@ 7 Ablation Studies
We conduct ablation studies to validate the contribution of our core components and design choices. Specifically, we examine four aspects: (1) the necessity of the two core loss terms, (2) the role of the projection head, (3) the impact of latent dimension, and (4) the effect of critical design choices on model performance.
7.1 Importance of Core Regularization Terms
Both slowness and uniformity losses are essential and complementary. We evaluate four model variants: a vanilla VAE, the full model, and two partial variants without $\mathcal{L}_{\text{uniform}}$ or $\mathcal{L}_{\text{slow}}$. For rollout evaluation, both partial variants diverged and produced NaN values, so we only report their loss statistics from autoencoder training. For completeness, we report the values of all loss terms during autoencoder training, even when they are not directly optimized. As shown in Table 2, removing either regularization leads to a substantial drop in rollout performance.
Table 2: Ablation study on the effect of regularization terms. We report the reconstruction loss $\mathcal{L}_{\text{recon}}$, the monitored slowness loss $\mathcal{L}_{\text{slow}}$, and the monitored uniformity loss $\mathcal{L}_{\text{uniform}}$.
Model $\mathcal{L}_{\text{recon}}$ $\mathcal{L}_{\text{slow}}$ $\mathcal{L}_{\text{uniform}}$
VAE-WM 0.00042 0.88 -3.04
GRWM 0.00067 0.11 -3.13
GRWM w/o $\mathcal{L}_{\text{uniform}}$ 0.00052 0.00015 0
GRWM w/o $\mathcal{L}_{\text{slow}}$ 0.00100 0.46 -2.47
When optimizing for slowness alone (w/o $\mathcal{L}_{\text{uniform}}$), we observe a classic case of representation collapse. The model aggressively minimizes the slowness loss by mapping all representations to a tiny region of the latent space, evidenced by a very high uniformity loss and low slowness loss. This confirms that $\mathcal{L}_{\text{uniform}}$ is indispensable for preventing trivial solutions.
When optimizing for uniformity alone (w/o $\mathcal{L}_{\text{slow}}$), we interestingly observe that the slowness metric naturally decreases. We attribute this effect to pushing different trajectories apart, which implicitly encourages representations from the same trajectory to cluster. However, explicitly including $\mathcal{L}_{\text{slow}}$ accelerates and reinforces this trend.
7.2 Role of the Projection Head
Table 3: Effect of the projection head. The projection head reduces reconstruction loss while maintaining latent probing performance.
Model $\mathcal{L}_{\text{recon}}$ Probing MSE
VAE-WM 0.00039 0.136
GRWM (w/ proj) 0.00061 0.058
GRWM (w/o proj) 0.00291 0.054
Figure 8: Frame-wise prediction MSE across ablation variants.
The projection head disentangles representation structuring from pixel-level reconstruction. While the model without a projection head can still learn a reasonably well-aligned latent space (as measured by latent probing MSE), its reconstruction loss is higher, as shown in table 3. By introducing the projection head, we allow the regularization losses to act in a separate subspace, freeing the primary latent space $z$ to focus on accurate reconstruction. This decoupling ultimately leads to better overall predictive performance and lower frame-wise MSE, as shown in figure 8.
7.3 Analysis on Latent Dimension
We conducted an ablation study on the dimensionality of the latent space, testing dimensions of 16, 32, 64, and 128. The results is presented in Figure 9.
Figure 9: Ablation study on the impact of latent dimension. GRWM (solid lines) consistently and significantly outperforms the vanilla VAE baseline (dashed lines) across all tested latent dimensions (16, 32, 64, and 128). Notably, our method’s performance is remarkably robust to the choice of latent dimension, while the baseline’s performance is highly sensitive.
The benefits of our regularization are independent of the latent space size. GRWM consistently outperforms the vanilla VAE-WM across all tested dimensions. For every capacity, the rollout error of GRWM (solid lines) is significantly lower than that of the corresponding baseline (dashed lines).
More importantly, our method demonstrates remarkable robustness to this hyperparameter. The performance curves for our model with latent dimensions 16, 32, 64, and 128 are nearly indistinguishable, indicating that our regularization technique successfully structures the latent space and learns a compact representation of the true state manifold, regardless of the available capacity. In contrast, the baseline’s performance is highly sensitive to the latent dimension. For the vanilla VAE-WM, a larger latent space appears to be detrimental, leading to faster error accumulation. Without proper regularization, a higher capacity latent space may overfit to irrelevant visual details or capture noise, which harms long-term prediction. Our method effectively mitigates this issue, ensuring stable and predictable performance, which is a highly desirable property for practical applications.
7.4 Design of the Slowness Loss
All-pairs temporal consistency is crucial for preventing degenerate solutions. A critical design choice in our ${\mathcal{L}}_{slow}$ formulation is to pull all pairs of frames within a trajectory’s context window closer, rather than only adjacent pairs. We compare our “All-Pairs” approach with an “Adjacent-Only” baseline. The results show that the “All-Pairs” strategy is superior. We attribute this to the causal nature of our encoder. Since the encoder’s output for frame $t$ is already conditioned on frames $t-k,\ldots,t-1$, simply minimizing the distance between $z_{t}$ and $z_{t-1}$ presents a “lazy” optimization problem due to their overlapping inputs. In contrast, our “All-Pairs” strategy enforces smoothness across distant frames with non-overlapping inputs (e.g., $z_{t}$ and $z_{t-k}$), leading to globally coherent latent trajectories and stronger long-horizon prediction.
@@ 8 Conclusion
Appendix
Appendix A Additional Rollout Visualizations
Appendix B Training Details
We summarize the training configurations for both the AutoEncoder and the dynamics models. Unless otherwise specified, models are trained with Adam optimizer and warmup-based learning rate schedules. The complete set of hyperparameters, including environment-specific settings, is provided in Table 4. For reproducibility, we will release the code and all configuration files upon paper acceptance.
Table 4: Training hyperparameters for AutoEncoder and Dynamics models.
AutoEncoder Training Setting
Epochs 50
Optimizer Adam (lr $5\times 10^{-4}$)
Scheduler Warmup-linear (1000 warmup, 10,000 total, min ratio 0.1)
Architecture Layers [2, 1, 2, 2, 1, 1, 2]; Encoder channels [256, 32]; Channels [256, 512, 512]
Maze $\times9$ $\lambda_{\text{uniform}}=\lambda_{\text{slow}}=0.1$, $\beta=1\times 10^{-6}$
Maze $\times3$ $\lambda_{\text{uniform}}=\lambda_{\text{slow}}=0.1$, $\beta=1\times 10^{-6}$
Minecraft $\lambda_{\text{uniform}}=\lambda_{\text{slow}}=0.01$, $\beta=1\times 10^{-6}$
Dynamics Models Training Setting
Optimizer Adam ($\beta_{1}=0.9,\ \beta_{2}=0.99$), weight decay $1\times 10^{-4}$
Scheduler Warmup-decay (1000 warmup, 100,000 total steps)
Diffusion 1000 steps, cosine $\beta$-schedule (shift=10.0), noise clip 20.0
Objective pred_v, fused-min-SNR loss weighting (clip=20.0, decay=0.9)
Sampling DDIM, 5 steps, $\eta=0.0$
VDM Classifier-free guidance, weight=5
Appendix C Dataset Details
@@
For the Maze environment, we fix the random seed to generate a consistent set of maps. We use Memory-Maze Environment (pasukonis2022memmaze). The rendered images are obtained using the MuJoCo engine. The agent has a discrete action space consisting of {move forward, turn left, turn right}. Trajectories are collected with a noisy A* algorithm to ensure sufficient coverage of the maze.
@@
For the Minecraft environment, we adopt the map from gornet2024automated. To restrict exploration, we enclose the area with wooden fences. Data collection is performed using the Malmo framework, with the same action space as in the Maze environment. The noisy A* policy is again used to generate diverse trajectories.

### v3

@@ 5.2 Temporal Contrastive Regularization
The output of the encoder, a sequence of latent vectors $\mathbf{z}\in\mathbb{R}^{B\times L\times\dots}$, is first passed through a projection layer (a linear layer) to produce embeddings $\mathbf{p}\in\mathbb{R}^{B\times L\times D}$ [^4] BB denotes the batch size, $L$ the sequence length, and $D$ the embedding dimension. [16]. These embeddings are then $L_{2}$-normalized to lie on the unit hypersphere: $\mathbf{p}^{\prime}=\frac{\mathbf{p}}{\|\mathbf{p}\|_{2}}$ [77]. Our regularization objectives are applied to these normalized embeddings, $\mathbf{p}^{\prime}$.
@@
The idea of temporal slowness is that consecutive or nearby states in a trajectory should have similar latent representations, reflecting the intuition that the underlying state of the environment evolves gradually over time [79]. Our loss encourages all pairs of frames within the same trajectory’s context window to be close to one another on the hypersphere. This enforces that the entire trajectory segment is mapped to a compact and continuous path in the representation space, ensuring that the latent representation evolves slowly and smoothly over time. We formalize this by minimizing the average L2 distance between all pairs of embeddings within a trajectory:
@@ 6.1 Setup
State-of-the-art latent generative world models generally adopt a standard VAE as the backbone for representation learning. Our method does not modify the dynamics model itself. Therefore, when comparing with existing sota approaches, we pair our representation module with the same dynamics models they use. We select three sota latent generative world models: Standard Diffusion (SD) [3], Video Diffusion Models (VD) [37], and Diffusion Forcing (DF)  [14]. A detailed description of training configurations is provided in the supplementary material.
@@ 6.3 Qualitative Long-Horizon Results
To further probe the long-term stability, we conduct an extreme long-horizon generation task. For this experiment, we specifically select sota world model, Diffusion Forcing, and the M9x9-DET and MC-DET dataset. In the following discussion, we refer to the original Diffusion Forcing as VAE-WM, and GRWM with Diffusion Forcing as GRWM. We use a sequence of randomly sampled actions to simulate an trajectory.
We evaluate our model on trajectories of varying lengths. To test long-horizon stability, we tasked the model with producing 10,000 consecutive frames from a single starting point. We visualize these long trajectories in Figure 4 and Figure 6. For the more complex MC-DET environment, we specifically evaluate a 100-frame rollout, as shown in Figure 5. Additional results under different starting conditions are provided in the supplementary material.
@@ 7 Ablation Studies
We conduct ablation studies to validate the contribution of our core components and design choices. Specifically, we examine four aspects: (1) the necessity of the two core loss terms, (2) the role of the projection head, (3) the impact of latent dimension, and (4) the effect of critical design choices on model performance. The detailed results are provided in the supplementary material.
8 Limitations
Environment Complexity.
The current evaluation is restricted to relatively simple and controlled environments. Scaling to more complex settings—such as partially observable, stochastic, or photorealistic environments—remains unresolved. These settings introduce additional structure and uncertainty, and require representations that can support consistent long-horizon predictions under such conditions.
Visual Artifacts.
GRWM still produces subtle visual artifacts and fails to recover fine-grained scene details, particularly in more complex environments such as Minecraft-DET. This indicates that the learned latent representation does not fully capture high-frequency or semantically relevant structure. Improving representation quality and decoding fidelity is a clear direction for future work.
Computation Overhead.
GRWM introduces additional computation beyond a standard VAE-based world model due to temporal aggregation and geometric regularization. Since the encoder uses a Transformer over temporal context, the overhead can increase as the sequence length grows.
Gap to Oracle Performance.
Although GRWM substantially narrows the gap to the oracle model, a large performance gap still remains. This suggests that even in deterministic environments, the learned representation is still far from sufficient to recover the full state information needed for near-perfect long-horizon prediction. Closing this gap remains an open problem, and further progress will likely require stronger representation learning methods.
@@ 9 Conclusion
Supplementary Material
Appendix A Training Details
Detailed training configurations, including hyperparameters and implementation details, are available in the accompanying code repository: https://github.com/XiaFire/Clone_Deterministic_Environment.
Appendix B Additional Evaluation Metrics
In deterministic environments, there exists a single ground-truth trajectory for a given initial condition and action sequence. As a result, pixel-wise MSE provides a direct and reliable measure of fidelity, and is used as our primary metric.
To complement this evaluation, we additionally report perceptual metrics on Maze $\times9$. These metrics capture structural similarity and distributional alignment that are not fully reflected by MSE. All metrics are computed in pixel space over rollout sequences. As shown in Table 2, the improvements are consistent with those observed under MSE.
Table 2: Supplementary perceptual metrics on Maze $\times9$.
SSIM $\uparrow$ rFID $\downarrow$
Model Baseline GR Baseline GR
[…]
Appendix C Additional Atari Experiments
We include additional experiments on Atari. The goal of this experiment is to extend our evaluation beyond static 3D environments to dynamic settings. We report PSNR and SSIM computed on rollout frames in pixel space. As shown in Table 3, GRWM consistently improves over the VAE-based world model across the evaluated environments. We note that this experiment is limited in scale and is not intended as a full Atari benchmark, but rather as a supplementary validation.
Table 3: Additional Atari results.
Env. PSNR $\uparrow$ SSIM $\uparrow$
VAE-WM GRWM VAE-WM GRWM
[…]
Appendix D Dataset Details
@@
For the Maze environment, we fix the random seed to generate a consistent set of maps. We use Memory-Maze Environment [57]. The rendered images are obtained using the MuJoCo engine. The agent has a discrete action space consisting of {move forward, turn left, turn right}. Trajectories are collected with a noisy A* algorithm to ensure sufficient coverage of the maze.
@@
For the Minecraft environment, we adopt the map from Gornet and Thomson [26] and enclose the area with wooden fences to restrict exploration. Trajectories are generated using a noisy A* policy under the same action space as in the Maze environment, providing diverse yet structured coverage. We first record the underlying deterministic state trajectories, and then render the corresponding pixel observations using Blender.
@@
Appendix E Ablation Studies
We conduct ablation studies to validate the contribution of our core components and design choices. Specifically, we examine four aspects: (1) the necessity of the two core loss terms, (2) the role of the projection head, (3) the impact of latent dimension, and (4) the effect of critical design choices on model performance. The detailed results are provided in the supplementary material.
Appendix F Additional Rollout Visualizations